OpenStudy (anonymous):

Find the probablitity of the given event: A single pair die is rolled. the number on the die is greater than 2

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

There are six faces on one die each with a different number: 1, 2, 3, 4, 5, or 6 The probability of any one number is 1/6. So, the probability that the number rolled is greater than 2 is the sum of the probabilities of getting a 3, 4, 5, or 6 = all of the options that are greater than 2. P(3)=1/6 P(4) = 1/6 p(5) = 1/6 P(6) = 1/6 P(>2) = 4/6 = 2/3